# Spin-spin entanglement in diffractive heavy quark production

### v1

@@ Abstract
Introduction—The recent experimental measurements of the spin-spin entanglement in top-antitop ($t\bar{t}$) quark pairs by the ATLAS Aad et al. (2024) and the CMS Hayrapetyan et al. (2024) collaborations at the LHC have garnered a significant attention amid a surge of interest in the intersection between collider physics and quantum information science Barr et al. (2024); Afik and de Nova (2022); Afik et al. (2025a). While quantum entanglement among elementary particles such as electrons and photons has been studied since the early days of quantum mechanics Einstein et al. (1935); Wu and Shaknov (1950); Bohm (1952); Bell (1964); Clauser et al. (1969), the direct observation of spin entanglement for quarks in QCD is highly nontrivial due to confinement. The top quarks are special because they decay by the electroweak interaction $t\to b+W^{+},\bar{t}\to\bar{b}+W^{-}$ before the strong interaction kicks in. The information about the spin state of the $t\bar{t}$ pair can be recovered by measuring the angular distribution of the leptons $\ell^{\pm}$ produced in the subsequent decays $W^{+}\to\ell^{+}+\nu_{\ell}$, $W^{-}\to\ell^{-}+\bar{\nu}_{\ell}$ Baumgart and Tweedie (2013); Afik and de Nova (2021); Aguilar-Saavedra (2023); Han et al. (2024).
@@
These quantum phenomena of course exist also in lighter quark systems such as bottom ($b\bar{b}$), charm ($c\bar{c}$) and strangeness ($s\bar{s}$) pairs. However, the corresponding measurements are more challenging because the pair first fragments hadronically. Fortunately, the spin state of the pair is partially retained when fragmenting into heavy baryons $q\to\Lambda_{q}$ and $\bar{q}\to\bar{\Lambda}_{q}$ Galanti et al. (2015) which then decay semi-leptonically. Although such processes suffer from small branching ratios, measurements may still be possible Kats and Uzan (2024); Afik et al. (2025b) given the high luminosity of the LHC (see also the recent $\Lambda\bar{\Lambda}$ measurement at RHIC STA (2025)).
@@
We assume the ‘Regge’ kinematics where the $\gamma^{*}+p$ center-of-mass (CM) energy $W^{2}=(p+q)^{2}$ is large. In the eikonal approximation, for the longitudinally polarized virtual photon, the cross section is given by Nikolaev et al. (2003); Altinoluk et al. (2016); Hatta et al. (2016); Boussarie et al. (2019).
$\hskip-17.07164pt\frac{d\sigma^{L}_{\alpha\alpha^{\prime}\beta\beta^{\prime}}}{dzd^{2}\boldsymbol{k}d^{2}{\bm{\Delta}}}=\frac{\alpha_{\rm em}e_{q}^{2}z\bar{z}Q^{2}}{N_{c}(q^{+})^{2}}\bigg{|}\int\frac{d^{2}\boldsymbol{p}T\left({\bm{p}}\right)}{(\boldsymbol{k}-\boldsymbol{p})^{2}+\mu^{2}}\bigg{|}^{2}$
@@
$\frac{d\sigma_{\alpha\alpha^{\prime}\beta\beta^{\prime}}^{L}}{dzd^{2}\boldsymbol{k}d^{2}{\bm{\Delta}}}=\frac{A^{L}}{4}\left(\delta_{\beta\alpha}\delta_{\beta^{\prime}\alpha^{\prime}}+C_{ab}^{L}\xi_{\beta}^{\dagger}\sigma^{a}\xi_{\alpha}\eta_{\beta^{\prime}}^{\dagger}\sigma^{b}\eta_{\alpha^{\prime}}\right),$ (8)
@@
Transverse photon—The calculation is significantly more complicated for the transversely polarized virtual photon. Again we write the $\gamma^{*}_{T}+p$ cross section in the form (8) with $L\to T$. The same cross section in the photo-production limit $Q=0$ is relevant to proton-nucleus UPCs Hagiwara et al. (2017); Reinke Pelicer et al. (2019); Gonçalves et al. (2020); Linek et al. (2023).
The unpolarized cross section is well known
@@
This takes the minimum value $C_{nn}^{T}=-1$ and $\Delta_{2}=2$ takes the maximal value if and only if $T_{1}=0$. (Naively, $C_{nn}^{T}$ also approaches 1-1 in the ultrarelativistic limit $k_{\perp}\to\infty$. However, $C_{nn}^{T}$ vanishes in this limit (cf., Fig. 1) due to a cancellation between $T_{1}$ and $T_{2}$.) Interestingly, in heavy quark production, $T_{1}$ does cross zero independently of models. To see this, assume that $k_{\perp}$ and $\mu$ are both much larger than the intrinsic transverse momentum $|{\bm{p}}|$ in (4). Then one can expand the denominator in $\bm{p}$}. Further using the color transparency condition $\int d^{2}{\bm{p}}T({\bm{p}})=0$ and neglecting the angular dependence of $T({\bm{p}})$, we find
@@
According to Gisin’s theorem Gisin (1991), for any two-qubit system that is a pure entangled state, the Bell-CHSH inequality is violated. Indeed, in the longitudinally polarized case, the density matrix satisfies the pure state condition $(\rho^{L})^{2}=\rho^{L}$. However, the transverse density matrix $\rho^{T}$ does not in general represents a pure state since $(\rho^{T})^{2}\neq\rho^{T}$. The equality holds only if $C_{nn}=-1$, in which case the state becomes pure and maximally entangled (22). This is depicted in Fig. 3 (bottom), to be compared with the nesting structure for generic two-qubit systems (top). The peculiarity of the present system is unmistakable.
@@
and their charge-conjugate counterparts for $\bar{b},\bar{c},\bar{s}$. The numbers above the first arrow are the fragmentation probabilities. The spin of the (anti)quark is largely inherited by the daughter (anti)baryon Galanti et al. (2015). The numbers above the second arrow are the branching ratios. Once these events have been identified and the baryon and antibaryon momenta have been reconstructed from their decay products Galanti et al. (2015); Kats and Uzan (2024); Afik et al. (2025b), one goes to their respective rest frames by first moving to the pair’s CM frame and then boosting along the $\pm\hat{k}$ directions. In each rest frame, one measures the angular distribution of one of the decay particles. In order to access the component $C_{nn}^{T}$, one needs to measure the polar angle of each decay particle with respect to the $\hat{n}$ axis in the form
$\frac{1}{\sigma}\frac{d\sigma^{\rm UPC}}{d\cos\theta_{+}d\cos\theta_{-}}=\frac{1+\alpha_{+}\alpha_{-}\cos\theta_{+}\cos\theta_{-}C_{nn}^{T}}{4},$ (25)
where $\alpha_{+}$ and $\alpha_{-}=-\alpha_{+}$ are the spin analyzing power of the measured particle and antiparticle. Equivalently,
$C_{nn}^{T}=\frac{9}{\alpha_{+}\alpha_{-}}\langle\cos\theta_{+}\cos\theta_{-}\rangle.$ (26)
Since $C_{nn}^{T}\leq 0$, we predict that $\langle\cos\theta_{+}\cos\theta_{-}\rangle\geq 0$, and this immediately signifies both entanglement and Bell-nonlocality. Maximal entanglement $C_{nn}^{T}=-1$ is achieved around $k_{\perp}\approx m$. To give a rough idea about the expected yield, we consider UPCs with the lead nucleus ($A=208,Z=82$) at the LHC energy $\sqrt{s_{NN}}=8.1$ TeV. According to Linek et al. (2023), $d\sigma_{c\bar{c}}^{\rm UPC}/dk_{\perp}\sim{\cal O}(100\,\mu{\rm b}/{\rm GeV})$ at $k_{\perp}\sim m_{c}$. We have confirmed this result and moreover predict that $d\sigma_{b\bar{b}}^{\rm UPC}/dk_{\perp}\sim{\cal O}(10\,{\rm nb}/{\rm GeV})$ at $k_{\perp}\sim m_{b}$ and $d\sigma_{s\bar{s}}^{\rm UPC}/dk_{\perp}\sim{\cal O}(100\,\mu{\rm b}/{\rm GeV})$. (Here we quote the value at $k_{\perp}\gtrsim 1$ GeV since, for the $s$-quark, the region $k_{\perp}\sim m_{s}$ is nonperturbative.) This suggests that, even after considering the suppression by the fragmentation and branching probabilities (24) (and other practical issues Kats and Uzan (2024); Afik et al. (2025b)), measurements may still be feasible thanks to the high luminosity of the LHC.
@@ Appendix A Spin density matrix element for the transversely polarized photon
$\frac{d\sigma^{T}_{\alpha\alpha^{\prime}\beta\beta^{\prime}}}{dzd^{2}\boldsymbol{k}d^{2}{\bm{\Delta}}}$ $=$ $\frac{\alpha_{em}e_{q}^{2}}{8z\bar{z}N_{c}(q^{+})^{2}}\int\frac{d^{2}\boldsymbol{p}T\left({\bm{p}}\right)}{(\boldsymbol{k}-\boldsymbol{p})^{2}+\mu^{2}}\int\frac{d^{2}\boldsymbol{p}^{\prime}T\left({\bm{p}}^{\prime}\right)}{(\boldsymbol{k}-\boldsymbol{p}^{\prime})^{2}+\mu^{2}}$ (27)
@@
$\xi_{\alpha}\xi^{\dagger}_{\beta}$ $=$ $\frac{1}{2}\left(\xi^{\dagger}_{\beta}\xi_{\alpha}+(\xi^{\dagger}_{\beta}\sigma^{n}\xi_{\alpha})\sigma^{n}+(\xi^{\dagger}_{\beta}\sigma^{r}\xi_{\alpha})\sigma^{r}+(\xi^{\dagger}_{\beta}\sigma^{k}\xi_{\alpha})\sigma^{k}\right)\;,$
$\tilde{\eta}_{-\beta^{\prime}}\tilde{\eta}^{\dagger}_{-\alpha^{\prime}}$ $=$ $\frac{1}{2}\left(\tilde{\eta}^{\dagger}_{-\alpha^{\prime}}\tilde{\eta}_{-\beta^{\prime}}+(\tilde{\eta}^{\dagger}_{-\alpha^{\prime}}\sigma^{n}\tilde{\eta}_{-\beta^{\prime}})\sigma^{n}+(\tilde{\eta}^{\dagger}_{-\alpha^{\prime}}\sigma^{r}\tilde{\eta}_{-\beta^{\prime}})\sigma^{r}+(\tilde{\eta}^{\dagger}_{-\alpha^{\prime}}\sigma^{k}\tilde{\eta}_{-\beta^{\prime}})\sigma^{k}\right)\;.$ (31)
Finally we use the formula $\tilde{\eta}_{-\alpha^{\prime}}^{\dagger}\vec{\sigma}\tilde{\eta}_{-\beta^{\prime}}=-\eta^{\dagger}_{\beta^{\prime}}\vec{\sigma}\eta_{\alpha^{\prime}}$ to recast the density matrix in the form (8). After tedious calculations, we find $C_{ab}^{T}=\widetilde{C}^{T}_{ab}/A^{T}$ where

### v2

@@ Abstract
Introduction—The recent experimental measurements of the spin-spin entanglement in top-antitop ($t\bar{t}$) quark pairs by the ATLAS Aad et al. (2024) and the CMS Hayrapetyan et al. (2024) collaborations at the LHC have garnered a significant attention amid a surge of interest in the intersection between collider physics and quantum information science Barr et al. (2024); Afik and de Nova (2022); Afik et al. (2025a). While quantum entanglement among elementary particles such as electrons and photons has been studied since the early days of quantum mechanics Einstein et al. (1935); Wu and Shaknov (1950); Bohm (1952); Bell (1964); Clauser et al. (1969), the direct observation of spin entanglement for quarks in QCD is highly nontrivial due to confinement. The top quarks are special because they decay by the electroweak interaction $t\to b+W^{+},\bar{t}\to\bar{b}+W^{-}$ before the strong interaction kicks in. The information about the spin state of the $t\bar{t}$ pair can be recovered by measuring the angular distribution of the leptons $\ell^{\pm}$ produced in the subsequent decays $W^{+}\to\ell^{+}+\nu_{\ell}$, $W^{-}\to\ell^{-}+\bar{\nu}_{\ell}$ Baumgart and Tweedie (2013); Afik and de Nova (2021); Aguilar-Saavedra (2023); Han et al. (2024); Dong et al. (2024).
@@
These quantum phenomena of course exist also in lighter quark systems such as bottom ($b\bar{b}$), charm ($c\bar{c}$) and strangeness ($s\bar{s}$) pairs. However, the corresponding measurements are more challenging because the pair first fragments hadronically. Fortunately, the spin state of the pair is partially retained when fragmenting into heavy baryons $q\to\Lambda_{q}$ and $\bar{q}\to\bar{\Lambda}_{q}$ Galanti et al. (2015); Lin et al. (2025) which then decay semi-leptonically. (See also an alternative approach Cheng and Yan (2025).) Although such processes suffer from small branching ratios, measurements may still be possible Kats and Uzan (2024); Afik et al. (2025b) given the high luminosity of the LHC. The recent $\Lambda\bar{\Lambda}$ measurement at RHIC STA (2025) is also encouraging.
@@
We assume the ‘Regge’ kinematics where the $\gamma^{*}+p$ center-of-mass (CM) energy $W^{2}=(p+q)^{2}$ is large. In the eikonal approximation, for the longitudinally polarized virtual photon, the cross section is given by Bartels et al. (1996); Nikolaev et al. (2003); Altinoluk et al. (2016); Hatta et al. (2016); Boussarie et al. (2019).
$\hskip-17.07164pt\left.\frac{d\sigma^{L}_{\alpha\alpha^{\prime}\beta\beta^{\prime}}}{dzd^{2}\boldsymbol{k}d^{2}{\bm{\Delta}}}\right|_{{\bm{\Delta}}\approx 0}=\frac{\alpha_{\rm em}e_{q}^{2}z\bar{z}Q^{2}}{N_{c}(q^{+})^{2}}\bigg|\int\frac{d^{2}\boldsymbol{p}T\left({\bm{p}}\right)}{(\boldsymbol{k}-\boldsymbol{p})^{2}+\mu^{2}}\bigg|^{2}$
@@
$\frac{d\sigma_{\alpha\alpha^{\prime}\beta\beta^{\prime}}^{L}}{dzd^{2}\boldsymbol{k}d^{2}{\bm{\Delta}}}=\frac{A^{L}}{4}\left(\delta_{\alpha\beta}\delta_{\alpha^{\prime}\beta^{\prime}}+C_{ab}^{L}\xi_{\alpha}^{\dagger}\sigma^{a}\xi_{\beta}\eta_{\alpha^{\prime}}^{\dagger}\sigma^{b}\eta_{\beta^{\prime}}\right),$ (8)
@@
Transverse photon—The calculation is significantly more complicated for the transversely polarized virtual photon. Again we write the $\gamma^{*}_{T}+p$ cross section in the form (8) with $L\to T$. The same cross section in the photo-production limit $Q=0$ is relevant to proton-nucleus UPCs Hagiwara et al. (2017); Reinke Pelicer et al. (2019); Gonçalves et al. (2020); Linek et al. (2023). The unpolarized cross section is well known
@@
This takes the minimum value $C_{nn}^{T}=-1$ and $\Delta_{2}=2$ takes the maximal value if and only if $T_{1}=0$. (Naively, $C_{nn}^{T}$ also approaches 1-1 in the ultrarelativistic limit $k_{\perp}\to\infty$. However, $C_{nn}^{T}$ vanishes in this limit (cf., Fig. 1) due to a cancellation between $T_{1}$ and $T_{2}$.) Interestingly, in heavy quark production, $T_{1}$ does cross zero. This can be shown model independently by expanding the integrand of (4) in $|{\bm{p}}|$. Using the color transparency condition $\int d^{2}{\bm{p}}T({\bm{p}})=0$, we find
@@
According to Gisin’s theorem Gisin (1991), for any two-qubit system that is a pure entangled state, the Bell-CHSH inequality is violated. Indeed, in the longitudinally polarized case, the density matrix satisfies the pure state condition $(\rho^{L})^{2}=\rho^{L}$. However, the transverse density matrix $\rho^{T}$ does not in general represents a pure state since $(\rho^{T})^{2}\neq\rho^{T}$. The equality holds only if $C_{nn}=-1$, in which case the state becomes pure and maximally entangled (22). This is depicted in Fig. 3 (bottom), to be compared with the nesting structure for generic two-qubit systems (top).
@@
and their charge-conjugate counterparts for $\bar{b},\bar{c},\bar{s}$. The numbers above the first arrow are the fragmentation probabilities. The spin of the (anti)quark is largely inherited by the daughter (anti)baryon Galanti et al. (2015). Similarly, especially for heavy quarks, roughly $z^{\rm quark}\approx z^{\rm hadron}$ and $k_{\perp}^{\rm quark}\approx k_{\perp}^{\rm hadron}$, to first approximation. The numbers above the second arrow are the branching ratios. Once these events have been identified and the baryon and antibaryon momenta have been reconstructed from their decay products Galanti et al. (2015); Kats and Uzan (2024); Afik et al. (2025b), one goes to their respective rest frames by first moving to the pair’s CM frame and then boosting along the $\pm\hat{k}$ directions. In each rest frame, the angular distribution of one of the decay particles can be measured. In order to access the component $C_{nn}^{T}$, one needs the polar angle distribution of each decay particle with respect to the $\hat{n}$ axis in the form
$\frac{1}{\sigma}\frac{d\sigma^{\rm UPC}}{d\cos\theta_{+}d\cos\theta_{-}}=\frac{1+\alpha_{+}\alpha_{-}r_{T}^{2}\cos\theta_{+}\cos\theta_{-}C_{nn}^{T}}{4},$ (25)
where $\alpha_{+}$ and $\alpha_{-}=-\alpha_{+}$ are the spin analyzing power of the measured particle and antiparticle, and $r_{T}$ is the transverse polarization retention factor Galanti et al. (2015). Equivalently,
$C_{nn}^{T}=\frac{9}{\alpha_{+}\alpha_{-}r_{T}^{2}}\langle\cos\theta_{+}\cos\theta_{-}\rangle.$ (26)
Since $C_{nn}^{T}\leq 0$, we predict that $\langle\cos\theta_{+}\cos\theta_{-}\rangle\geq 0$, and this signifies both entanglement and Bell-nonlocality, although the effect of hadronization should be carefully investigated. To give a rough idea about the expected yield, we consider UPCs with the lead nucleus ($A=208,Z=82$) at the LHC energy $\sqrt{s_{NN}}=8.1$ TeV. According to Linek et al. (2023), $d\sigma_{c\bar{c}}^{\rm UPC}/dk_{\perp}\sim{\cal O}(100\,\mu{\rm b}/{\rm GeV})$ at $k_{\perp}\sim m_{c}$. We have confirmed this result and moreover predict that $d\sigma_{b\bar{b}}^{\rm UPC}/dk_{\perp}\sim{\cal O}(10\,{\rm nb}/{\rm GeV})$ at $k_{\perp}\sim m_{b}$ and $d\sigma_{s\bar{s}}^{\rm UPC}/dk_{\perp}\sim{\cal O}(100\,\mu{\rm b}/{\rm GeV})$. (Here we quote the value at $k_{\perp}\gtrsim 1$ GeV since, for the $s$-quark, the region $k_{\perp}\sim m_{s}$ is nonperturbative.) The fragmentation and branching probabilities (24) bring in a suppression factor of order $10^{-4}$ for the $s\bar{s}$ pair and $10^{-6}$ for the $c\bar{c},b\bar{b}$ pairs. Further considering $\alpha_{\pm},r_{T}\sim{\cal O}(1)$ Galanti et al. (2015), we find the signal cross sections of order 10nb, 100pb, 0.01fb for the $s\bar{s},c\bar{c},b\bar{b}$ channels, respectively. Assuming $1\sim 10$pb-1 of proton-lead collision data at the future high-luminosity LHC experiment d’Enterria et al. (2025), our rough estimates indicate that the $s\bar{s}$-channel is feasible, while the $c\bar{c}$-channel is marginal and the $b\bar{b}$-channel is impractical at the LHC. In future, the discussion will be extended to light quarks Hatta and Schoenleber (2025) that have larger cross sections. Also, other methods to measure the spin correlation without relying on weak decays should be explored Cheng and Yan (2025).
@@ Appendix A Spin density matrix element for the transversely polarized photon
$\left.\frac{d\sigma^{T}_{\alpha\alpha^{\prime}\beta\beta^{\prime}}}{dzd^{2}\boldsymbol{k}d^{2}{\bm{\Delta}}}\right|_{{\bm{\Delta}}\approx 0}$ $=$ $\frac{\alpha_{em}e_{q}^{2}}{8z\bar{z}N_{c}(q^{+})^{2}}\int\frac{d^{2}\boldsymbol{p}T\left({\bm{p}}\right)}{(\boldsymbol{k}-\boldsymbol{p})^{2}+\mu^{2}}\int\frac{d^{2}\boldsymbol{p}^{\prime}T\left({\bm{p}}^{\prime}\right)}{(\boldsymbol{k}-\boldsymbol{p}^{\prime})^{2}+\mu^{2}}$ (27)
@@
$\xi_{\beta}\xi^{\dagger}_{\alpha}$ $=$ $\frac{1}{2}\left(\xi^{\dagger}_{\alpha}\xi_{\beta}+(\xi^{\dagger}_{\alpha}\sigma^{n}\xi_{\beta})\sigma^{n}+(\xi^{\dagger}_{\alpha}\sigma^{r}\xi_{\beta})\sigma^{r}+(\xi^{\dagger}_{\alpha}\sigma^{k}\xi_{\beta})\sigma^{k}\right)\;,$
$\tilde{\eta}_{-\alpha^{\prime}}\tilde{\eta}^{\dagger}_{-\beta^{\prime}}$ $=$ $\frac{1}{2}\left(\tilde{\eta}^{\dagger}_{-\beta^{\prime}}\tilde{\eta}_{-\alpha^{\prime}}+(\tilde{\eta}^{\dagger}_{-\beta^{\prime}}\sigma^{n}\tilde{\eta}_{-\alpha^{\prime}})\sigma^{n}+(\tilde{\eta}^{\dagger}_{-\beta^{\prime}}\sigma^{r}\tilde{\eta}_{-\alpha^{\prime}})\sigma^{r}+(\tilde{\eta}^{\dagger}_{-\beta^{\prime}}\sigma^{k}\tilde{\eta}_{-\alpha^{\prime}})\sigma^{k}\right)\;.$ (31)
Finally we use the formula $\tilde{\eta}_{-\beta^{\prime}}^{\dagger}\vec{\sigma}\tilde{\eta}_{-\alpha^{\prime}}=-\eta^{\dagger}_{\alpha^{\prime}}\vec{\sigma}\eta_{\beta^{\prime}}$ to recast the density matrix in the form (8). After tedious calculations, we find $C_{ab}^{T}=\widetilde{C}^{T}_{ab}/A^{T}$ where